OpenStudy (anonymous):

What is the solution of this equation? WORK SHOWN! 22=r-18

OpenStudy (texaschic101):

22 = r - 18 ---- add 18 to both sides 22 + 18 = r - 18 + 18 --- whatever you do on one side of the equal sign, you have to do on the other. Notice how the 18's cancel out on the right side. 40 = r any questions ?

OpenStudy (reni_love_ela):